A ball of mass 1.0kg is moving in a circle of radius 1.0m. Calculate its acceleration if it moves with a constant angular speed of 2.0 rads-1. Suppose the ball starts from rest and moves with a constant angular acceleration of 0.02 rads-2. What is the net force on the ball 60 s after it starts moving from rest ?


1
Expert's answer
2020-09-30T05:01:15-0400
  • mass of ball(m)=0.1kg

radius of circle (r)=1.0m

angular velocity ("\omega)" =2.0 rads-1

angular acceleration of ball("\alpha" )-

"\alpha" = r"\omega^2"

= (1.0)(22)

= 4 rads-2

  • We need to calculate angular velocity of vall after 60 sec, so

"\alpha" = "\frac{\omega_1-\omega_2}{t}"

0.02 = "\frac{\omega_1-0}{60}"

"\omega_1" = 1.2 rads-1

Net Force on ball(F) = mr"\omega_1^2"

= (0.1)(1.0)(1.22)

= 0.144 N